For the reaction

X(g)+4Y(g)⇌3Z(g)

Kp=3.30×10−2 bar−2 at a temperature of 375 ∘C .

Calculate the value of Kc.
Answer #1

아e: i barı =0996923 am 2. 乙 kp- 3.3 10 am we mow thah 2 ke = 3.37 -2乂2830 .3 = 95-66 フ